OpenStudy (anonymous):

solve the diffrential equation : sqrt(1-x^2) dy/dx=y+1 , y(0) = 1

OpenStudy (anonymous):

\[\sqrt{1-x ^{2}}\frac{dy}{dx}=y+1\]\[\frac{1}{y+1}dy=\frac{1}{\sqrt{1-x ^{2}}}dx\]\[\int\limits_{}^{} \frac{1}{y+1}dy=\int\limits_{}^{} \frac{1}{\sqrt{1-x ^{2}}}dx\]

OpenStudy (anonymous):

tanx